# Understanding Generalizability of Diffusion Models Requires Rethinking the Hidden Gaussian Structure

### v1

@@ 3.2 Inductive Bias towards Learning the Gaussian Structures
•  Similarity in weight matrices. As illustrated in Figure 4(left), $\bm{W}_{\sigma(t)}$ progressively converge towards $\bm{U}\tilde{\Lambda}_{\sigma(t)}\bm{U}^{T}$ throughout the linear distillation process, achieving small normalized MSE (less than 0.2) for most of the noise levels. The less satisfactory convergence behavior at $\sigma(t)=80.0$ is due to inadequate training of the diffusion models at this particular noise level, which is minimally sampled during the training of actual diffusion models (see Appendix G for more details).
@@ 4.2 Overparameterized Models Learn Gaussian Structures before Memorization
Diffusion models learn the Gaussian structures with generalizability before it memorizes.
Figure 8(a) demonstrates that during early training epochs (0-841), $\mathcal{D}_{\bm{\theta}}$ progressively converge to $\mathcal{D}_{\mathrm{G}}$ in the intermediate noise variance regime, indicating that the diffusion model is progressively learning the Gaussian structure in the initial stages of training. Notably. By epoch 841, the diffusion model generates images strongly resembling those produced by the Gaussian model, as shown in Figure 8(b). However, continued training beyond this point increases the score approximation error as the model transitions toward memorization. This observation suggests that early stopping could be an effective strategy for promoting generalization in overparameterized diffusion models.
@@ 5 Connection between Strong Generalizability and Gaussian Structure
Comparing Figure 9(a)(top) and (b)(top), we observe that $\mathcal{D}_{\mathrm{G}}$ generates nearly identical images whether the Gaussian structure is calculated on a small dataset (1,094 images) or a much larger one (70,000 images). This similarity emerges because datasets of the same class can exhibit similar Gaussian structure (empirical covariance) with relatively few samples—just hundreds for FFHQ. Given the Gaussian structure’s critical role in generalization, small datasets may already contain much of the information for generalization, contrasting previous assertions in [20] that strong generalization requires training on datasets of substantial size (more than $10^{5}$ images). However, smaller datasets increase memorization risk, as shown in Figure 9(b). To mitigate this, as discussed in Section 4, we can either reduce model capacity or implement early stopping (Figure 9(c)). Indeed, models trained on 1094 and 35000 images generate remarkably similar images, though the smaller dataset yields lower perceptual quality. This similarity further demonstrates that small datasets contain substantial generalization-relevant information closely tied to Gaussian structure.
@@ F.1 Behaviors in Low-noise Regime
We visualize the score differences between $\mathcal{D}_{\mathrm{G}}$ and $\mathcal{D}_{\bm{\theta}}$ in low-noise regime in Figure 16. The left figure demonstrates that when the dataset size becomes smaller than a certain threshold, the score difference at $\sigma=0$ remains persistently non-zero. Moreover, the right figure shows that this difference depends solely on dataset size rather than model capacity. This phenomenon arises from two key factors: (i) $\mathcal{D}_{\bm{\theta}}$ approximates an identity mapping at low noise levels, independent of training dataset size, and (ii) $\mathcal{D}_{\mathrm{G}}$ equals an identity mapping only when the empirical covariance matrix is low-rank.
@@ G.5 Measuring Score Approximation Error with NMSE
Appendix H Computing Resources
@@
All the diffusion models in the experiments are trained on A100 GPUs provide by NCSA Delta GPU [33].

### v2

@@ 3.2 Inductive Bias towards Learning the Gaussian Structures
•  Similarity in weight matrices. As illustrated in Figure 4(left), $\bm{W}_{\sigma(t)}$ progressively converge towards $\bm{U}\tilde{\Lambda}_{\sigma(t)}\bm{U}^{T}$ throughout the linear distillation process, achieving small normalized MSE (less than 0.2) for most of the noise levels. The less satisfactory convergence behavior at $\sigma(t)=80.0$ is due to inadequate training of the diffusion models at this particular noise level, which is minimally sampled during the training of actual diffusion models (see Section G.2 for more details).
@@ 4.2 Overparameterized Models Learn Gaussian Structures before Memorization
Diffusion models learn the Gaussian structures with generalizability before they memorize.
Figure 8(a) demonstrates that during early training epochs (0-841), $\mathcal{D}_{\bm{\theta}}$ progressively converge to $\mathcal{D}_{\mathrm{G}}$ in the intermediate noise variance regime, indicating that the diffusion model is progressively learning the Gaussian structure in the initial stages of training. Notably. By epoch 841, the diffusion model generates images strongly resembling those produced by the Gaussian model, as shown in Figure 8(b). However, continued training beyond this point increases the difference between $\mathcal{D}_{\bm{\theta}}$ and $\mathcal{D}_{\mathrm{G}}$ as the model transitions toward memorization. This observation suggests that early stopping could be an effective strategy for promoting generalization in overparameterized diffusion models.
@@ 5 Connection between Strong Generalizability and Gaussian Structure
Comparing Figure 9(a)(top) and (b)(top), we observe that $\mathcal{D}_{\mathrm{G}}$ generates nearly identical images whether the Gaussian structure is calculated on a small dataset (1,094 images) or a much larger one (70,000 images). This similarity emerges because datasets of the same class can exhibit similar Gaussian structure (empirical covariance) with relatively few samples—just hundreds for FFHQ. Given the Gaussian structure’s critical role in generalization, small datasets may already contain much of the information for generalization, contrasting previous assertions in [20] that strong generalization requires training on datasets of substantial size (more than $10^{5}$ images). However, smaller datasets increase memorization risk, as shown in Figure 9(b). To mitigate this, as discussed in Section 4, we can either reduce model capacity or implement early stopping (Figure 9(c)). Indeed, models trained on 1094 and 35000 images generate remarkably similar images, though the smaller dataset yields lower perceptual quality. This similarity further demonstrates that small datasets contain substantial generalization-relevant information closely tied to Gaussian structure. Further discussion on the connections and differences between our work and [20] are detailed in Appendix H.
@@ F.1 Behaviors in Low-noise Regime
We visualize the score differences between $\mathcal{D}_{\mathrm{G}}$ and $\mathcal{D}_{\bm{\theta}}$ in low-noise regime in Figure 16. The left figure demonstrates that when the dataset size becomes smaller than a certain threshold, the score difference at $\sigma=0$ remains persistently non-zero. Moreover, the right figure shows that this difference depends solely on dataset size rather than model capacity. This phenomenon arises from two key factors: (i) $\mathcal{D}_{\bm{\theta}}$ converges to the identity mapping at low noise levels, independent of training dataset size and model capacity, and (ii) $\mathcal{D}_{\mathrm{G}}$ approximates the identity mapping at low noise levels only when the empirical covariance matrix is full-rank, as can be seen from (6).
@@ G.5 Measuring Score Approximation Error with NMSE
Appendix H Discussion on Geometry-Adaptive Harmonic Bases
H.1 GAHB only Partially Explain the Strong Generalization
Recent work [20] observes that diffusion models trained on sufficiently large non-overlapping datasets (of the same class) generate nearly identical images. They explain this ”strong generalization” phenomenon by analyzing bias-free deep diffusion denoisers with piecewise linear input-output mappings:
$\mathcal{D}(\bm{x}_{t};\sigma(t))$ $=\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))\bm{x}$ (54)
$=\sum_{k}\lambda_{k}(\bm{x}_{t})\bm{u}_{k}(\bm{x}_{t})\bm{v}_{k}^{T}(\bm{x}_{t})\bm{x}_{t},$ (55)
where $\lambda_{k}(\bm{x}_{t})$, $\bm{u}_{k}(\bm{x}_{t})$, and $\bm{v}_{k}(\bm{x}_{t})$ represent the input-dependent singular values, left and right singular vectors of the network Jacobian $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$. Under this framework, strong generalization occurs when two denoisers $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ have similar Jacobians: $\nabla\mathcal{D}_{1}(\bm{x}_{t};\sigma(t))\approx\nabla\mathcal{D}_{2}(\bm{x}_{t};\sigma(t))$. The authors conjecture this similarity arises from networks’ inductive bias towards learning certain optimal $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ that has sparse singular values and the singular vectors of which are the geometry-adaptive harmonic bases (GAHB)—near-optimal denoising bases that adapt to input $\bm{x}_{t}$.
While [20] provides valuable insights, their bias-free assumption does not reflect real-world diffusion models, which inherently contain bias terms. For feed forward ReLU networks, the denoisers are piecewise affine:
$\mathcal{D}(\bm{x}_{t};\sigma(t))$ $=\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))\bm{x}_{t}+\bm{b}_{\bm{x}_{t}},$ (56)
where $\bm{b}_{\bm{x}_{t}}$ is the network bias that depends on both network parameterization and the noisy input $\bm{x}_{t}$ [44]. Here, similar Jacobians alone cannot explain strong generalization, as networks may differ significantly in $\bm{b}_{\bm{x}_{t}}$. For more complex network architectures where even piecewise affinity fails, we consider the local linear expansion of $\mathcal{D}(\bm{x}_{t};\sigma(t))$:
$\mathcal{D}(\bm{x}_{t}+\Delta\bm{x};\sigma(t))$ $=\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))\Delta\bm{x}_{t}+\mathcal{D}(\bm{x}_{t};\sigma(t)),$ (57)
which approximately holds for small perturbation $\Delta\bm{x}$. Thus, although $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ characterizes $\mathcal{D}(\bm{x}_{t};\sigma(t))$’s local behavior around $\bm{x}_{t}$, it does not provide sufficient information on the global properties.
Our work instead examines global behavior, demonstrating that $\mathcal{D}(\bm{x}_{t};\sigma(t))$ is close to $\mathcal{D}_{\mathrm{G}}(\bm{x}_{t};\sigma(t))$—the optimal linear denoiser under the Gaussian data assumption. This implies that strong generalization partially stems from networks learning similar Gaussian structures across non-overlapping datasets of the same class. Since our linear model captures global properties but not local characteristics, it complements the local analysis in [20].
H.2 GAHB Emerge only in Intermediate-Noise Regime
For completeness, we study the evolution of the Jacobian matrix $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across various noise levels $\sigma(t)$. The results are presented in Figures 28 and 29, which reveal three distinct regimes:
  • High-noise regime [10,80]. In this regime, the leading singular vectors [^6] We only care about leading singular vectors since the Jacobians in this regime are highly low-rank. The less well aligned singular vectors have singular values near 0. of the Jacobian matrix $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ well align with those of the Gaussian weights (the leading principle components of the training dataset), consistent with our finding that diffusion denoisers approximate linear Gaussian denoisers in this regime. Notice that DAEs trained sufficiently on separate noise levels (Figure 29) show stronger alignment compared to vanilla diffusion models (Figure 28), which suffer from insufficient training at high noise levels.
  • Intermediate-noise regime [0.1,10]: In this regime, GAHB emerge as singular vectors of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ diverge from the principal components, becoming increasingly adaptive to the geometry of input image.
  • Low-noise regime [0.002,0.1]. In this regime, the leading singular vectors of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ show no clear patterns, consistent with our observation that diffusion denoisers approach the identical mapping, which has unconstrained singular vectors.
Notice that the leading singular vectors of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ are the input directions that lead to the maximum variation in denoised outputs, thus revealing meaningful information on the local properties of $\mathcal{D}(\bm{x}_{t};\sigma(t))$ at $\bm{x}_{t}$. As demonstrated in Figure 30, perturbing input $\bm{x}_{t}$ along these vectors at difference noise regimes leads to distinct effects on the final generated images: (i) in the high-noise regime where the leading singular vectors align with the principle components of the training dataset, perturbing $\bm{x}_{t}$ along these directions leads to canonical changes such as image class, (ii) in the intermediate-noise regime where the GAHB emerge, perturbing $\bm{x}_{t}$ along the leading singular vectors modify image details such as colors while preserving overall image structure and (iii) in the low-noise regime where the leading singular vectors have no significant pattern, perturbing $\bm{x}_{t}$ along these directions yield no meaningful semantic changes.
These results collectively demonstrate that the singular vectors of the network Jacobian $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ have distinct properties at different noise regimes, with GAHB emerging specifically in the intermediate regime. This characterization has significant implications for uncertainty quantification [45] and image editing [46].
Figure 28: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels. Figure(a) shows the generation trajectory. Figure(b) shows the correlation matrix between Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ and training dataset principal components $\bm{U}$. Notice that the leading singular vectors of $\bm{U}(\bm{x}_{t})$ and $\bm{U}$ well align in early timesteps but diverge in later timesteps. Figure(c) shows the first three principle components of the training dataset while figures(d-f) show the evolution of Jacobian’s first three singular vectors across noise levels. These singular vectors initially match the principle components but progressively adapt to input image geometry, before losing distinct patterns at very low noise levels. While we present only left singular vectors, right singular vectors exhibit nearly identical behavior and yield equivalent results.
Figure 29: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels for DAEs. We repeat the experiments in Figure 28 on DAEs that are sufficiently trained on each discrete noise levels. Notice that with sufficient training, the Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ show a better alignment with principle components $\bm{U}$ in early timesteps.
Figure 30: Effects of perturbing $\bm{x}_{t}$ along Jacobian singular vectors. Figure(a)-(c) demonstrate the effects of perturbing input $\bm{x}_{t}$ along the first singular vector of the Jacobian matrix ($\bm{x}_{t}\pm\lambda\bm{u}_{1}(\bm{x}_{t})$) on the final generated images. Perturbing $\bm{x}_{t}$ in high-noise regime (Figure (a)) leads to canonical image changes while perturbation in intermediate-noise regime (Figure (b)) leads to change in details but the overall image structure is preserved. At very low noise variances, perturbation has no significant effect (Figure (c)). Similar effects are observed in concurrent work [46].
Appendix I Computing Resources
@@
All the diffusion models in the experiments are trained on A100 GPUs provided by NCSA Delta GPU [33].